OpenStudy (anonymous):

Shannon's scooter has a piece of tape stuck to the tire. If the tire has a diameter of 18 inches, how far does the piece of tape travel in 72° of rotation?

OpenStudy (anonymous):

dont you find the radians of the degree first?

OpenStudy (anonymous):

This is a pretty interesting and a fun yet simple question! it's just all about how well you imagine this in your head :)

OpenStudy (anonymous):

Okay so lets draw ourselves a circle first, shall we?  our \[d = 18 \]

OpenStudy (anonymous):

if our d = 18 then our r = 9 right? okay you should start imagining the arc in this, if our degree is 72° (which is a positive degree, we can consider this positive degree moving counterclockwise) like this: 

OpenStudy (anonymous):

now you can convert the degree to radians if you wish, to convert 72° to degree to radians then you should probably be familiar with the pi and 180° ratio \[\pi rad = 180°\] \[\frac{ \pi }{ 180 }rad = 1°\]

OpenStudy (anonymous):

so if 1° gives \[\frac{ \pi }{ 180 }rad\] then 72° should give \[\frac{ 2 }{ 5 }\pi rad\]

OpenStudy (anonymous):

the formula to find the arc length in a radian is \[S = \theta*r\]

OpenStudy (anonymous):

we have our radius and we have our theta so all you need to do is plug the numbers into the formula \[S = \frac{ 2\pi }{ 5 } * 9\] \[S = \frac{ 18 }{ 5 }\pi\] so the tape moved 11.309 inches